# Learning to Pursue AC Optimal Power Flow Solutions with Feasibility Guarantees

### v1

@@ Abstract
This paper focuses on an AC optimal power flow (OPF) problem for distribution feeders equipped with controllable distributed energy resources (DERs). We consider a solution method that is based on a continuous approximation of the projected gradient flow – referred to as the safe gradient flow – that incorporates voltage and current information obtained either through real-time measurements or power flow computations. These two setups enable both online and offline implementations. The safe gradient flow involves the solution of convex quadratic programs (QPs). To enhance computational efficiency, we propose a novel framework that employs a neural network approximation of the optimal solution map of the QP. The resulting method has two key features: (a) it ensures that the DERs’ setpoints are practically feasible, even for an online implementation or when an offline algorithm has an early termination; (b) it ensures convergence to a neighborhood of a strict local optimizer of the AC OPF. The proposed method is tested on a 93-node distribution system with realistic loads and renewable generation. The test shows that our method successfully regulates voltages within limits during periods with high renewable generation.
@@ I Introduction
Contributions. In this paper, we consider a solution method for the AC OPF that is based on a continuous approximation of the projected gradient flow – hereafter referred to as the safe gradient flow [23, 24] – incorporating voltage and current information obtained either through real-time measurements (as in feedback optimization [2, 3, 4, 5]) or power flow computations. To favor computational efficiency and speed, we propose a novel framework that employs a neural network approximation of the safe gradient flow. In particular, the neural network predicts the unique optimal solution of a quadratic program (QP) defining the map of the safe gradient flow. The learning task is well-posed, in the sense that the optimal solution map of the QP is a single-valued function and it is continuous. The learned safe gradient flow is then used in conjunction with voltage and current information to identify AC OPF solutions. We summarize our contributions as follows:
(c1) We propose an iterative method where the neural network approximation of the safe gradient flow is used with either real-time measurements or power flow computations.
(c2) We show that our method leads to solutions that are practically feasible. The term practical feasibility refers to the fact that we provide guarantees on the maximum constraint violation (which is found to be negligible through our numerical experiments); the analytical estimate of the violation allows for a careful tightening of the constraints in the AC OPF so that the neural network can be trained not to violate the actual constraints. The practical feasibility is at any time, in the sense that the algorithm produces feasible points even when terminated before convergence or implemented online.
@@
(c4) We perform numerical experiments on a 93-bus distribution system [25] and with realistic load and solar production profiles from the Open Power System Data. We show that our approach ensures voltage regulation and satisfaction of the DERs’ constraints. Our method shows far superior performance in terms of voltage regulation compared to approaches that attempt to approximate the solutions of the OPF directly.
The remainder of the paper is organized as follows. Section II will formulate the AC OPF and will explain our proposed mathematical model. Section III will provide details on the neural network-based safe gradient flow, while Section IV will illustrate simulation results. Section V will present our theoretical results, and Section VI will conclude the paper.
@@ II Problem Formulation and Proposed Model
II-A Distribution System Model
We consider a distribution system [^1] Notation. We use the following notational conventions throughout the paper. Boldface upper-case letters (e.g., $\mathbf{X}$) denote matrices, and boldface lower-case letters (e.g., $\mathbf{x}$) denote column vectors. The transpose of a vector or matrix is denoted by $(\cdot)^{\top}$, and the complex conjugate by $(\cdot)^{*}$. The imaginary unit is denoted by $j$, satisfying $j^{2}=-1$, and the absolute value of a scalar is written as ||$\cdot$|| ⋅ |. For a real-valued vector $\mathbf{x}\in\mathbb{R}^{N}$, $\mathrm{diag}(\mathbf{x})$ returns an $N\times N$ diagonal matrix with the entries of $\mathbf{x}$ on the diagonal. The $\ell_{2}$-norm of a vector $\mathbf{x}\in\mathbb{R}^{n}$ is denoted $\|\mathbf{x}\|$; for a matrix $\mathbf{X}\in\mathbb{R}^{n\times m}$, $\|\mathbf{x}\|$ is the induced $\ell_{2}$-norm. For two vectors $\mathbf{x}\in\mathbb{R}^{n}$ and $\mathbf{u}\in\mathbb{R}^{m}$, the notation $(\mathbf{x},\mathbf{u})\in\mathbb{R}^{n+m}$ denotes their concatenation. The symbol $\mathbf{0}$ is used to denote vectors or matrices of zeros, with dimension determined from context. The set of complex numbers is denoted $\mathbb{C}$. For a complex vector $\boldsymbol{x}\in\mathbb{C}^{N}$, $\Re(\boldsymbol{x})\in\mathbb{R}^{N}$ denotes its real part and $\Im(\boldsymbol{x})\in\mathbb{R}^{N}$ its imaginary part. We denote by $\mathbb{N}_{0}$ the set of non-negative integers, and by $\mathbb{N}_{>0}$ the set of positive integers. The set of all integers is denoted by $\mathbb{Z}$. comprising $N+1$ nodes, labeled by $\{0,1,\dots,N\}$. Node 00 represents the substation (or point of common coupling), whereas $\mathcal{N}:=\{1,\dots,N\}$ contains the remaining nodes; these nodes may feature a mix of uncontrollable loads and controllable DERs. We focus on a steady-state representation in which currents and voltages are modeled as complex phasors. For each node $k\in\mathcal{N}$, let the line-to-ground voltage phasor be $v_{k}=\nu_{k}e^{j\,\delta_{k}}\in\mathbb{C},$ with magnitude $\nu_{k}=|v_{k}|$ and angle $\delta_{k}$. The phasorial representation of the current injected at node $k$ is $i_{k}=|i_{k}|e^{j\,\psi_{k}}\in\mathbb{C}$. At the substation, the voltage is denoted as $v_{0}=V_{0}e^{j\,\delta_{0}}$ [26].
As usual, applying Ohm’s and Kirchhoff’s Laws in the phasor domain yields the relationship
$\begin{bmatrix}i_{0}\\[3.0pt] \boldsymbol{i}\end{bmatrix}\;=\;\begin{bmatrix}y_{0}&\boldsymbol{\bar{y}}^{\top}\\[3.0pt] \boldsymbol{\bar{y}}&\boldsymbol{Y}\end{bmatrix}\begin{bmatrix}v_{0}\\[3.0pt] \boldsymbol{v}\end{bmatrix},$ (1)
where $\boldsymbol{i}=[\,i_{1},\dots,i_{N}\,]^{\mathsf{T}}\!\in\mathbb{C}^{N}$ and $\boldsymbol{v}=[\,v_{1},\dots,v_{N}\,]^{\mathsf{T}}\!\in\mathbb{C}^{N}$, and where the admittance matrix $\boldsymbol{Y}\!\in\!\mathbb{C}^{N\times N}$ and the vectors $\boldsymbol{\bar{y}}\!\in\!\mathbb{C}^{N},y_{0}\in\mathbb{C}$ are built based on the series and shunt parameters of the lines under a $\Pi$-model [26].
@@ II-A Distribution System Model
where $\boldsymbol{s}_{l}:=\boldsymbol{p}_{l}+j\boldsymbol{q}_{l}\in\mathbb{C}^{N}$ is a vector collecting the aggregate complex powers of non-controllable loads at each node. Finally, we consider a set of nodes $\mathcal{M}\subseteq\mathcal{N}$ where voltages are monitored, and let $M=|\mathcal{M}|$ denote the number of such nodes.
Given the controllable powers $\boldsymbol{u}$ and the loads $\boldsymbol{p}_{l},\boldsymbol{q}_{l}$, one can employ numerical techniques to solve (2) for the voltages $\boldsymbol{v}$. It is important to note that the power flow equation (2) may admit zero, one, or multiple solutions [27, 28, 29]. If multiple solutions exist, we focus on practical solutions; i.e., the solution within the neighborhood of the nominal voltage profile that yields relatively high voltage magnitudes and low line currents. Due to the Implicit Function Theorem, we can define a map $(\boldsymbol{u},\boldsymbol{s}_{l})\mapsto\boldsymbol{v}(\boldsymbol{u},\boldsymbol{s}_{l})$ mapping loads and power from the DERs into complex voltages at the nodes. Additionally, based on the function $\boldsymbol{v}(\boldsymbol{u},\boldsymbol{s}_{l})$ and the topology of the network, we also define the function $(\boldsymbol{u},\boldsymbol{s}_{l})\mapsto\boldsymbol{i}(\boldsymbol{u},\boldsymbol{s}_{l})$ mapping loads and power from the DERs into line currents; we assume that we monitor a set $\mathcal{E}$ of $L$ lines.
The functions $\boldsymbol{v}(\boldsymbol{u},\boldsymbol{s}_{l})$ and $\boldsymbol{i}(\boldsymbol{u},\boldsymbol{s}_{l})$ are utilized to formulate instances of the AC OPF. In the remainder of this paper, we proceed under the following practical assumption.
Assumption 1 (Maps in a neighborhood of the nominal voltage profile).
@@
This assumption is supported by the findings in, e.g., [27, 29, 28]. This assumption will be utilized only in the analysis of the algorithms; it will not play a role in the algorithmic design and practical implementations of the proposed methods.
Remark II.1 (Model and notation).
We note that the framework proposed in this paper is applicable to multi-phase distribution systems with both wye and delta connections under the same Assumption 1. However, to simplify the notation and to streamline the exposition, we outline the framework using a single-phase model. $\Box$
II-B AC OPF Formulation
Several formulations for the AC OPF at the distribution level has been proposed in the literature; see, for example, the survey [1] and the representative works [30, 31, 3, 5]. In this section, we structure our presentation around an AC OPF formulation that includes constraints on node voltages, line currents, and operating ranges of DERs.
Recall that $\boldsymbol{u}=[p_{1},\dots,p_{G},q_{1},\dots,q_{G}]^{\top}\in\mathbb{R}^{2G}$ represents the vector of DER active and reactive power injections, and recall that $\mathcal{M}\subseteq\mathcal{N}$ is a set of nodes where voltages are monitored and controlled. In particular, for the latter, let lower and upper bounds on the voltage magnitudes be denotes as $\underline{V}$ and $\overline{V}$, respectively. Additionally, let $\overline{I}$ be an ampacity limit for the $L$ lines that are monitored. We then consider the following problem formulation to compute the DERs’ power setpoints:
$\textsf{U}^{*}(\boldsymbol{s}_{l},\boldsymbol{\theta}):=\arg\min_{\boldsymbol{u}\in\mathcal{C}}\quad$ $C_{v}(\boldsymbol{v}(\boldsymbol{u};\boldsymbol{s}_{l}))+C_{p}(\boldsymbol{u})$
s.t. $\underline{V}\leq|\boldsymbol{v}(\boldsymbol{u};\boldsymbol{s}_{l})|\leq\overline{V},$
$|\boldsymbol{i}(\boldsymbol{u};\boldsymbol{s}_{l})|\leq\overline{I},$ (3)
@@ II-B AC OPF Formulation
where $C_{v}:\mathbb{R}^{M}\rightarrow\mathbb{R}$ is a cost associated with the voltage profile, $C_{p}:\mathbb{R}^{2G}\rightarrow\mathbb{R}$ captures DER-specific costs, and the set $\mathcal{C}_{i}(\boldsymbol{\theta}_{u,i})$ encodes constraints for the $i$th DER, such as capacity and hardware limits as well as grid code requirements. The inequalities in the voltage and current constraints are taken entry-wise. We allow a parametric representation of the set through parameters $\boldsymbol{\theta}_{u,i}$; to this end, we assume the set $\mathcal{C}_{i}(\boldsymbol{\theta}_{u,i})$ can be expressed as
@@
where $\ell_{i}$ is a vector-valued function modeling power limits, and the inequality is taken entry-wise. The function $\ell_{i}$ is assumed to be differentiable. For example, if the $i$th DER is an inverter-interfaced controllable renewable source, then $\ell_{i}(p_{i},q_{i},\boldsymbol{\theta}_{u,i})=[p_{i}^{2}+q_{i}^{2}-s_{n,i}^{2},p_{i}-p_{\text{max},i},-p_{i}]^{\top}$, where $\boldsymbol{\theta}_{u,i}=(p_{\text{max},i},s_{n,i})$ with $s_{n,i}$ and $p_{\text{max},i}$ the inverter rated size and the maximum available active power, respectively. The overall set of inputs that parametrize the problem (II-B) is denoted as $\boldsymbol{\theta}:=(\boldsymbol{\theta}_{u,i},\dots,\boldsymbol{\theta}_{u,G},\underline{V},\bar{V},\bar{I})$; these inputs are in addition to $\boldsymbol{s}_{l}$. We will use the notation $\mathcal{C}=\mathcal{C}_{1}\times\ldots\times\mathcal{C}_{G}$ and we define the set $\mathcal{S}(\boldsymbol{s}_{l}):=\mathcal{S}_{v}(\boldsymbol{s}_{l})\cap\mathcal{S}_{i}(\boldsymbol{s}_{l})$, where:
$\mathcal{S}_{v}(\boldsymbol{s}_{l})$ :=$:=\{\boldsymbol{u}\in\mathcal{C}:\underline{V}\leq|\boldsymbol{v}(\boldsymbol{u};\boldsymbol{s}_{l})|\leq\overline{V}\}$
$\mathcal{S}_{i}(\boldsymbol{s}_{l})$ :=$:=\{\boldsymbol{u}\in\mathcal{C}:|\boldsymbol{i}(\boldsymbol{u};\boldsymbol{s}_{l})|\leq\overline{I}\}\,.$
The feasible set of (II-B) is $\mathcal{S}_{v}(\boldsymbol{s}_{l})\cap\mathcal{S}_{i}(\boldsymbol{s}_{l})\cap\mathcal{C}$. In the following, for notational simplicity, we drop the dependence on $\boldsymbol{s}_{l}$.
It is well known that the AC OPF is nonconvex and may admit multiple globally optimal and locally optimal solutions. Accordingly, the function $(\boldsymbol{s}_{l},\boldsymbol{\theta})\mapsto\textsf{U}^{*}(\boldsymbol{s}_{l},\boldsymbol{\theta})$ that maps parameters of the problem into globally optimal solutions to the AC OPF is a set-valued function. Since identifying a solution $\boldsymbol{u}^{*}\in\textsf{U}^{*}(\boldsymbol{s}_{l},\boldsymbol{\theta})$ is in general difficult, we consider the (sub-)set of $\textsf{U}^{*}(\boldsymbol{s}_{l},\boldsymbol{\theta})$ that contains points that are local minimizers and isolated KKT points for (II-B); we denote such set as $\textsf{U}^{\textsf{lm}}(\boldsymbol{s}_{l},\boldsymbol{\theta})$ (although we note that some local minimizers can also be global minimizers). In the following, we explain our approach to identify local minimizers in $\textsf{U}^{\textsf{lm}}(\boldsymbol{s}_{l},\boldsymbol{\theta})$.
II-C Proposed Mathematical Framework and Implementations

### v2

@@ Abstract
This paper focuses on an AC optimal power flow (OPF) problem for distribution feeders equipped with controllable distributed energy resources (DERs). We consider a solution method that is based on a continuous approximation of the projected gradient flow – referred to as the safe gradient flow – that incorporates voltage and current information obtained either through real-time measurements or power flow computations. These two setups enable both feedback-based online and offline implementations. The safe gradient flow involves the solution of convex quadratic programs (QPs). To enhance computational efficiency, we propose a novel framework that employs a neural network approximation of the optimal solution map of the QP. The resulting method has two key features: (a) it ensures that the DERs’ setpoints are practically feasible, even for an online implementation or when an offline algorithm has an early termination; (b) it ensures convergence to a neighborhood of a strict local optimizer of the AC OPF. The proposed method is tested on a 93-node distribution system with realistic loads and renewable generation. The test shows that our method successfully regulates voltages within limits during periods with high renewable generation.
@@ I Introduction
Contributions. In this paper, we consider a solution method for the AC OPF that is based on the so called safe gradient flow [1, 11], a constrained gradient flow whose design employs Control Barrier Functions (CBFs), modified to incorporate voltage and current information obtained either through real-time measurements (as in feedback optimization [19, 12, 5, 33]) or power flow computations. To favor computational efficiency and speed, we propose a novel framework that employs a neural network approximation of the safe gradient flow. In particular, the neural network predicts the unique optimal solution of a quadratic program (QP) defining the map of the safe gradient flow. The learning task is well-posed, in the sense that the optimal solution map of the QP is a single-valued function and it is continuous. The learned safe gradient flow is then used in conjunction with voltage and current information to identify AC OPF solutions. We summarize our contributions as follows:
(c1) We propose an iterative method where the neural network approximation of the safe gradient flow is used with either real-time measurements or power flow computations. When using real-time measurements, the proposed method constitutes a neural network-based feedback optimization method.
(c2) We show that our method leads to solutions that are practically feasible. The term practical feasibility refers to the fact that we provide guarantees on the maximum constraint violation; the analytical estimate of the violation allows for a careful tightening of the constraints in the AC OPF so that the neural network can be trained not to violate the actual constraints. The practical feasibility is at any time, in the sense that the algorithm produces feasible points even when terminated before convergence or implemented online. To the best of our knowledge, this is the first neural network-based algorithm for the AC OPF problem with feasibility guarantees, without postprocessing.
@@
(c4) Numerical simulations are performed on a 93-bus distribution grid [35] with realistic time-series for load and renewable generation from the Open Power System Data. The considered network [35] is a low voltage benchmark distribution system, reflecting realistic operational settings. We show that our approach ensures voltage regulation and satisfaction of the DERs’ constraints. Our proposed method shows superior performance in terms of voltage regulation compared to approaches that attempt to approximate the solutions of the OPF directly, even when using more data to train the latter.
Finally, we note that this paper provides several innovations relative to our previous work [11]. We propose a novel framework using neural network approximations of the safe gradient flow to solve the AC optimal power flow problem; the proposed method leads to implementations for both online optimization and offline computation, providing new computationally-efficient methods for both applications (as shown shortly in Figure 1). We provide results for convergence and feasibility that are new in the literature, and we test the neural network-based solution on a voltage regulation problem. We provide tighter bounds when we characterize the forward invariance of an inflated set of feasible points. These improved results can be used to effectively train the proposed NN-SGF under tightened constraints to ensure forward invariance of the original feasible set.
The remainder of the paper is organized as follows. Section II formulates the AC OPF and explains our proposed mathematical model. Section III provides details on the neural network-based safe gradient flow, while Section IV illustrates simulation results. Section V presents our theoretical results and Section VI gathers our conclusions. Finally, the appendix contains the technical proofs. [^1] Notation. We use the following notational conventions throughout the paper. Boldface upper-case letters (e.g., $\mathbf{X}$) denote matrices, and boldface lower-case letters (e.g., $\mathbf{x}$) denote column vectors. The transpose of a vector or matrix is denoted by $(\cdot)^{\top}$, and the complex conjugate by $(\cdot)^{*}$. The imaginary unit is denoted by $j$, satisfying $j^{2}=-1$, and the absolute value of a scalar is written as $|\cdot|$. For notational simplicity, given a vector-valued function $\mathbf{f}(\mathbf{x}):\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$, the notation $|\mathbf{f}(\mathbf{x})|$ means that the absolute value is taken entry-wise; i.e., $|\mathbf{f}(\mathbf{x})|=[|f_{1}(\mathbf{x})|,\ldots,|f_{m}(\mathbf{x})|]^{\top}$. For a real-valued vector $\mathbf{x}\in\mathbb{R}^{N}$, $\mathrm{diag}(\mathbf{x})$ returns an $N\times N$ diagonal matrix with the entries of $\mathbf{x}$ on the diagonal. The $\ell_{2}$-norm of a vector $\mathbf{x}\in\mathbb{R}^{n}$ is denoted $\|\mathbf{x}\|$; for a matrix $\mathbf{X}\in\mathbb{R}^{n\times m}$, $\|\mathbf{X}\|$ is the induced $\ell_{2}$-norm. The $\ell_{1}$-norm of a vector $\mathbf{x}\in\mathbb{R}^{n}$ is denoted $\|\mathbf{x}\|_{1}$, and the $\ell_{\infty}$-norm is denoted $\|\mathbf{x}\|_{\infty}$. For two vectors $\mathbf{x}\in\mathbb{R}^{n}$ and $\mathbf{u}\in\mathbb{R}^{m}$, the notation $(\mathbf{x},\mathbf{u})\in\mathbb{R}^{n+m}$ denotes their concatenation. The symbol $\mathbf{0}$ is used to denote vectors or matrices of zeros, with dimension determined from context. The set of complex numbers is denoted $\mathbb{C}$. For a complex vector $\boldsymbol{x}\in\mathbb{C}^{N}$, $\Re(\boldsymbol{x})\in\mathbb{R}^{N}$ denotes its real part and $\Im(\boldsymbol{x})\in\mathbb{R}^{N}$ its imaginary part. We denote by $\mathbb{N}_{0}$ the set of non-negative integers, and by $\mathbb{N}_{>0}$ the set of positive integers. The set of all integers is denoted by $\mathbb{Z}$.
@@ II Problem Formulation and Proposed Model
The paper considers a distribution system with $N+1$ nodes, labeled by $\{0,1,\dots,N\}$. Node 0 represents the substation (or point of common coupling), whereas $\mathcal{N}:=\{1,\dots,N\}$ contains the remaining nodes; these nodes may feature a mix of uncontrollable loads and controllable DERs. We focus on a steady-state representation in which currents and voltages are modeled as complex phasors. For each node $k\in\mathcal{N}$, let the line-to-ground voltage phasor be $v_{k}\in\mathbb{C}$. The current injected at node $k$ is $i_{k}$. At the substation, the voltage is denoted as $v_{0}=V_{0}e^{j\,\delta_{0}}$ [24]. As usual, applying Ohm’s and Kirchhoff’s Laws in the phasor domain yields the relationship
$\begin{bmatrix}i_{0}\\[3.0pt] \boldsymbol{i}_{\mathcal{N}}\end{bmatrix}\;=\;\begin{bmatrix}y_{0}&\boldsymbol{\bar{y}}^{\top}\\[3.0pt] \boldsymbol{\bar{y}}&\boldsymbol{Y}\end{bmatrix}\begin{bmatrix}v_{0}\\[3.0pt] \boldsymbol{v}_{\mathcal{N}}\end{bmatrix}$, (1)
where $\boldsymbol{i}_{\mathcal{N}}=[\,i_{1},\dots,i_{N}\,]^{\mathsf{T}}\!\in\mathbb{C}^{N}$ and $\boldsymbol{v}_{\mathcal{N}}=[\,v_{1},\dots,v_{N}\,]^{\mathsf{T}}\!\in\mathbb{C}^{N}$, and where the admittance matrix $\boldsymbol{Y}\!\in\!\mathbb{C}^{N\times N}$ and the vectors $\boldsymbol{\bar{y}}\!\in\!\mathbb{C}^{N},y_{0}\in\mathbb{C}$ are built based on the series and shunt parameters of the lines under a $\Pi$-model [24].
@@
where $\boldsymbol{p}_{l},\boldsymbol{q}_{l}\in\mathbb{R}^{N}$ are the vectors collecting the aggregate active and reactive powers, respectively, of non-controllable loads at each node. Hereafter, we let $\boldsymbol{s}_{l}:=[\boldsymbol{p}_{l}^{\top},\boldsymbol{q}_{l}^{\top}]^{\top}$. Finally, nodes $\mathcal{M}\subseteq\mathcal{N}$, $M=|\mathcal{M}|$, are monitored and voltages can be measured, and lines $\mathcal{L}=\{1,\ldots,L\}\subset\mathcal{N}\times\mathcal{N}$ are also monitored and currents can be measured.
In this formulation, we consider a standard setup where node 0 is designated as the reference bus with $V_{0}=1$ p.u. and $\delta_{0}=0$. All remaining nodes $k\in\mathcal{N}$ are treated as PQ buses. Inverter-interfaced distributed energy resources (DERs) are represented as controllable active and reactive power injections $(p_{i},q_{i})$ at their respective interconnection nodes. Power injections are defined on a net basis, accounting for all elements connected at each node: injections are positive at nodes hosting DERs, negative at pure load nodes, and net values at nodes hosting both generation and demand. Given the controllable powers $\boldsymbol{u}$ and the loads $\boldsymbol{p}_{l},\boldsymbol{q}_{l}$, one can employ numerical techniques to solve (2) for the voltages $\boldsymbol{v}$ (for example, the Newton–Raphson algorithm or fixed-point methods) [10, 42]. It is important to note that the power flow equation (2) may admit zero, one, or multiple solutions [8, 6, 42]. If multiple solutions exist, we focus on practical solutions; i.e., the solution within the neighborhood of the nominal voltage profile that yields relatively high voltage magnitudes and low line currents. Due to the Implicit Function Theorem, we can define a map $(\boldsymbol{u},\boldsymbol{s}_{l})\mapsto\boldsymbol{v}(\boldsymbol{u},\boldsymbol{s}_{l}):=[v_{1}(\boldsymbol{u},\boldsymbol{s}_{l}),\ldots,v_{M}(\boldsymbol{u},\boldsymbol{s}_{l})]^{\top}$, mapping loads and power from the DERs into voltages at the monitored nodes. Additionally, based on the function $\boldsymbol{v}(\boldsymbol{u},\boldsymbol{s}_{l})$ and the topology of the network, we also define the function $(\boldsymbol{u},\boldsymbol{s}_{l})\mapsto\boldsymbol{i}(\boldsymbol{u},\boldsymbol{s}_{l}):=[i_{1}(\boldsymbol{u},\boldsymbol{s}_{l}),\ldots,i_{L}(\boldsymbol{u},\boldsymbol{s}_{l})]^{\top}$, mapping loads and DERs’ powers into monitored line currents.
The functions $\boldsymbol{v}(\boldsymbol{u},\boldsymbol{s}_{l})$ and $\boldsymbol{i}(\boldsymbol{u},\boldsymbol{s}_{l})$ are utilized to formulate instances of the AC OPF. We adopt the notation $|\boldsymbol{v}(\boldsymbol{u},\boldsymbol{s}_{l})|$ and $|\boldsymbol{i}(\boldsymbol{u},\boldsymbol{s}_{l})|$ to denote the entry-wise absolute value of these vector-valued functions; i.e., $|\boldsymbol{v}(\boldsymbol{u},\boldsymbol{s}_{l})|=[|v_{1}(\boldsymbol{u},\boldsymbol{s}_{l})|,\dots,|v_{M}(\boldsymbol{u},\boldsymbol{s}_{l})|]^{\top}$ and $|\boldsymbol{i}(\boldsymbol{u},\boldsymbol{s}_{l})|=[|i_{1}(\boldsymbol{u},\boldsymbol{s}_{l})|,\dots,|i_{L}(\boldsymbol{u},\boldsymbol{s}_{l})|]^{\top}$. In the remainder of this paper, we proceed under the following practical assumption.
Assumption 1 (Neighborhood of nominal voltage).
@@
This assumption is supported by the findings in, e.g., [8, 42, 6]. This assumption will be utilized only in the analysis of the algorithms; it will not play a role in the algorithmic design and implementations of the proposed methods.
II-A AC OPF Formulation
Several formulations for the AC OPF at the distribution level have been proposed in the literature; see, for example, the survey [27] and the representative works [7, 18, 12, 33]. In this section, we structure our presentation around an AC OPF formulation that includes constraints on node voltages, line currents, and operating ranges of DERs.
Recall that $\boldsymbol{u}=[p_{1},\dots,p_{G},q_{1},\dots,q_{G}]^{\top}\in\mathbb{R}^{2G}$ represents the vector of DER active and reactive power injections, and recall that $\mathcal{M}\subseteq\mathcal{N}$ is a set of nodes where voltages are monitored and controlled. In particular, for the latter, let lower and upper bounds on the voltage magnitudes be denoted as $\underline{V}$ and $\overline{V}$, respectively. Additionally, let $\overline{I}$ be an ampacity limit for the lines that are monitored. We then consider the following problem formulation to compute the DERs’ power setpoints:
$\textsf{U}^{*}(\boldsymbol{s}_{l},\boldsymbol{\theta}):=\arg\min_{\boldsymbol{u}\in\mathcal{C}}\quad$ $\phi_{v}(|\boldsymbol{v}(\boldsymbol{u};\boldsymbol{s}_{l})|)+\phi_{p}(\boldsymbol{u})$
s.t. $\underline{V}\leq|v_{j}(\boldsymbol{u};\boldsymbol{s}_{l})|\leq\overline{V},~~\forall j\in\mathcal{M}$
$\|i_{j}(\boldsymbol{u};\boldsymbol{s}_{l})|\leq\overline{I},~~~~~~\forall j\in\mathcal{L}$ (3)
@@ II-A AC OPF Formulation
where $\phi_{v}:\mathbb{R}^{M}\rightarrow\mathbb{R}$, $\phi_{p}:\mathbb{R}^{2G}\rightarrow\mathbb{R}$, the set $\mathcal{C}_{i}(\boldsymbol{\theta}_{u,i})$ encodes constraints for the $i$th DER, and the inequalities in the voltage and current constraints are taken entry-wise. The cost $\phi_{p}(\boldsymbol{u})$ models DER-related considerations such as minimizing active-power curtailment of renewable resources, limiting excessive reactive power provision to reduce inverter losses, or incorporating economic incentives and fairness criteria among DERs. The voltage-related cost $\phi_{v}(|\boldsymbol{v}|)$ serves as a performance objective and complements the hard voltage constraints. While the constraints enforce admissible voltage limits, $\phi_{v}(|\boldsymbol{v}|)$ allows the operator to penalize deviations from a desired voltage profile (e.g., nominal voltages) or promote flatter voltage profiles across the network.
We allow a parametric representation of the set through parameters $\boldsymbol{\theta}_{u,i}$; to this end, we assume the set $\mathcal{C}_{i}(\boldsymbol{\theta}_{u,i})$ can be expressed as
@@
where $\ell_{i}$ is a vector-valued function modeling power limits, and the inequality is taken entry-wise. The function $\ell_{i}$ is assumed to be differentiable. For example, if the $i$th DER is an inverter-interfaced controllable renewable source, then $\ell_{i}(p_{i},q_{i},\boldsymbol{\theta}_{u,i})=[p_{i}^{2}+q_{i}^{2}-s_{n,i}^{2},p_{i}-p_{\text{max},i},-p_{i}]^{\top}$, where $\boldsymbol{\theta}_{u,i}=(p_{\text{max},i},s_{n,i})$ with $s_{n,i}$ and $p_{\text{max},i}$ the inverter rated size and the maximum available active power, respectively. The overall set of inputs that parametrize the problem (II-A) is denoted as $\boldsymbol{\theta}:=(\boldsymbol{\theta}_{u,i},\dots,\boldsymbol{\theta}_{u,G},\underline{V},\bar{V},\bar{I})$; these inputs are in addition to $\boldsymbol{s}_{l}$. We will use the notation $\mathcal{C}=\mathcal{C}_{1}\times\ldots\times\mathcal{C}_{G}$ and we define the set $\mathcal{S}(\boldsymbol{s}_{l}):=\mathcal{S}_{v}(\boldsymbol{s}_{l})\cap\mathcal{S}_{i}(\boldsymbol{s}_{l})$, where:
$\mathcal{S}_{v}(\boldsymbol{s}_{l})$ $:=\{\boldsymbol{u}\in\mathcal{C}:\underline{V}\leq|v_{j}(\boldsymbol{u};\boldsymbol{s}_{l})|\leq\overline{V},j\in\mathcal{M}\}$
$\mathcal{S}_{i}(\boldsymbol{s}_{l})$ $:=\{\boldsymbol{u}\in\mathcal{C}:|i_{j}(\boldsymbol{u};\boldsymbol{s}_{l})|\leq\overline{I},j\in\mathcal{L}\}\,.$
The feasible set of (II-A) is $\mathcal{S}_{v}(\boldsymbol{s}_{l})\cap\mathcal{S}_{i}(\boldsymbol{s}_{l})$. In the following, for notational simplicity, we drop the dependence on $\boldsymbol{s}_{l}$.
It is well known that the AC OPF is nonconvex and may admit multiple globally optimal and locally optimal solutions. Accordingly, the function $(\boldsymbol{s}_{l},\boldsymbol{\theta})\mapsto\textsf{U}^{*}(\boldsymbol{s}_{l},\boldsymbol{\theta})$ that maps parameters of the problem into globally optimal solutions to the AC OPF is a set-valued function. Since identifying a solution $\boldsymbol{u}^{*}\in\textsf{U}^{*}(\boldsymbol{s}_{l},\boldsymbol{\theta})$ is in general difficult, we consider a set of local minimizers and isolated KKT points for (II-A); we denote such set as $\textsf{U}^{\textsf{lm}}(\boldsymbol{s}_{l},\boldsymbol{\theta})$ (although we note that some local minimizers can also be global minimizers). In the following, we explain our approach to identify local minimizers in $\textsf{U}^{\textsf{lm}}(\boldsymbol{s}_{l},\boldsymbol{\theta})$.
II-B Proposed Mathematical Framework and Implementations